Find the areas under the normal curve in each of the following cases: (2pts. each)



1. Between z = 0 and z = 1.63 2. Between z = 0 and tau = 0.98



3. To the right of z= 0.29



4. To the left of



5. Between z=-1.56 and z=-1.83



z = - 1.39

1
Expert's answer
2022-05-31T08:44:37-0400

1.


"P(0<Z<1.63) = P(Z<1.63) \u2013 P(Z<0)"

"=0.948449-0.5=0.448449"

2.


"P(0<Z\le0.98) = P(Z\le0.98) – P(Z\le0)"

"=0.836457-0.5=0.336457"

3.


"P(Z>0.29) = 1-P(Z\le0.29)"

"=0.385908"

4.


"P(Z<-1.39) =0.082264"



5.


"P(-1.83<Z<-1.56) = P(Z<-1.56)"

"– P(Z\le-1.83)=0.0593799-0.0336250"

"=0.025755"